OpenStudy (anonymous):

Find the relative extrema for the function. f(x,y)=4x/x^2+y^2+1

OpenStudy (anonymous):

@stamp

OpenStudy (anonymous):

already looked at that. I'm looking for a partial derivative check because its messy.

OpenStudy (stamp):

\[f(x,y)=\frac{4x}{x^2+y^2+1}\]

OpenStudy (anonymous):

The second partial gets messy*

OpenStudy (stamp):

ok so I solved partial with respect to x on paper and it checks with wolfram. Not sure how to find the critical point though. Do we need the second order partial?\[f_x=\frac{-4x^2+4y^2+4}{(x^2+y^2+1)^2}\]

OpenStudy (anonymous):

yeah I think I got the 2nd partial finally because I re-wrote the equation so I could use product rule instead of quotient rule.

OpenStudy (stamp):

I am working on solving fxx

OpenStudy (anonymous):

\[\frac{ (-4x^2+4y^2+1)(-4x) }{ (x^2+y^2+1)^3 }-\frac{ 8x }{ (x^2+y^2+1)^2}\]

OpenStudy (stamp):

I have been trying to work this out on paper and It has been destroying me, how did you rearrange it for the product rule?

OpenStudy (anonymous):

yeah the same thing happened to me when I tried it the first time. I set \[f_x=(-4x^2+4y^2+4)*(x^2+y^2+1)^-2\]

OpenStudy (anonymous):

my final answer also has an error but its very minor

OpenStudy (stamp):

\[f_{xx}=\frac{8x(x^2-3y^2-3)}{(x^2+y^2+1)^3}\]So what do we do with it

OpenStudy (stamp):

do we have to find\[f_y,\ f_{yy},\ f_{xy}\]

OpenStudy (anonymous):

I will type out what i got

OpenStudy (anonymous):

\[f_y=\frac{ -8xy }{ (x^2+y^2+1)^2 }\]

OpenStudy (anonymous):

\[f_\left( yy \right)= \frac{ 32xy^2 }{ (x^2+y^2+1)^3 }-\frac{ 8y }{ (x^2+y^2+1)^2 }\]

OpenStudy (anonymous):

\[f_ \left( yx \right)=\frac{ 32x^2y }{ (x^2+y^2+1)^3 }-\frac{ 8y }{ (x^2+y^2+1)^2 }\]

OpenStudy (anonymous):

my yy should say -8x

OpenStudy (stamp):

I simplified my fxx and fyy to share a denominator, yours are separated. but ok we have fx, fxx, fy, fyy. What now

OpenStudy (stamp):

set a system to solve fx and fy = 0

OpenStudy (anonymous):

This is the part where I get confused if you go to the lamar notes they explain how to find the relative min and maxes but I'm not sure if I'm doing it right

OpenStudy (anonymous):

your first link now makes a bit of sense to me now

OpenStudy (stamp):

well all our partials , when simplified into one fraction, have the same denominators so we just set the numerators equal and solve for the numerators = 0

OpenStudy (stamp):

check my links to see the partials simplified

OpenStudy (stamp):

man i have been at work this whole time i will check this out later when i have free time